OpenStudy (chosenmatt):

Identify which property of equality is used to transform Equation 1 to Equation 2. Equation 1: x – 6 = 14 Equation 2: x = 20 A. multiplication property of equality B. subtraction property of equality C. division property of equality D. addition property of equality

OpenStudy (texaschic101):

x - 6 = 14 --- add 6 to both sides x - 6 + 6 = 14 + 6 x = 20 addition property of equality

OpenStudy (chosenmatt):

u sure it's not multiplication?

OpenStudy (anonymous):

It's Subtraction!!!!